OpenStudy (anonymous):

Given (-5,1) is on the graph of f(x) ..What's the corresponding point for f(5x)?

OpenStudy (anonymous):

Is it? (1,-1)

OpenStudy (anonymous):

is this a multiple choice question? We can use the equation: \[\huge y-y_o=m(x-x_o)\] where \[\huge x_o=-5 \space \space and \space \space y_0=1\]

OpenStudy (anonymous):

x_o = x "not" or "knot" as in, the first x term. does that make sense (in case you didn't know)